OpenStudy (anonymous):

4 times the measure of the complement of an angle is equal to / the measure of the original angle's supplement. What is the measure of the angle?

OpenStudy (anonymous):

Recall that the complement of an angle x is 90-x, and the supplement of an angle is 180-x. So you can convert this statement into an equation as follows: 4(90-x)=(180-x) which you can then solve for x.

OpenStudy (anonymous):

i dont know what to do with the equation

OpenStudy (anonymous):

Ok, so for the left side you can distribute the four first. You then get 360-4x=180-x Can you see where we go with this equation?

OpenStudy (anonymous):

640-4x=180

OpenStudy (anonymous):

is that right